Two particles A and B are released from rest, having equal charges of -8x10^-6 C and initially separated from each other by 0.025 m. The masses of particle A and B are 6.25x10^-6 kg and 3.75x10^-6 kg respectively. Find the magnitude and initial accelerations of particle A and B.


1
Expert's answer
2022-04-04T14:43:14-0400

"ma=\frac{kq^2}{r^2},"

"a=\frac{kq^2}{mr^2},"

"a_A=0.15\cdot 10^9~\frac m{s^2},"

"a_B=0.25\cdot 10^9~\frac m{s^2}."